You go on a round trip in a very fast spacecraft,

non-calc based Physics
Answer #1

Use the equation for Time dilation

t'= t/ Sqrt(1-v2/c2)

25= 5/Sqrt(1-v2/c2)

625=25/1-v2/c2

1-v2/c2=25/625

v2/c2= 1-25/625

v2/c2= 600/625

v/c=0.9798

v=0.9798c